

********************************************************************
UARS WEEKLY STATUS REPORT - 11 MAY 2001
********************************************************************
***********
SOLSTICE
***********
The cause of the operational anomaly last week was determined to be a wrong
load sent to the instrument. No damage occured because of this error.
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
The calibration of SUSIM using three of its deuterium lamps begun last
week was successfully completed. Our thanks go to the UARS Project
Engineers, the Flight Operations, and Mission Planning Teams for
planning and directing the UARS spacecraft and ground systems for
closed loop SSPP pointing.
After the calibration, SUSIM performed its daily solar scans for the
remainder of the week during periods of predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM/AXIS was activated in its full power mode on April 30 and was
deactivated on May 7. At the current time, AXIS is not generating energy
deposition. SwRI is looking into this. Tests of AXIS production code on
the rehost computer are continuing.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
PEM has examined its processed files of the rehost machine and has found
that no differences exist in the Level 1 and Level 2 output files which were
generated the rehost when the same input data was used on CDHF and CDHFB.
PEM is now confident that output generated by PEM 3.32, Jobs 1 & 2, on the
rehost system is valid and correct.
The PEM team is working on replacing the PEM:: machine with a DEC Alpha
machine to be called PEM2K::. Tests of the NRT data transfer from CDHF2::
through PEM2K:: to the unix computer running the PEM GSE software show no
difference between equivalent NRT files transferred from PEM::. Thus, NRT
data from CDHF2:: is captured correctly on PEM2K::. Tests of the NRT data
transfer from CDHFB:: were not successful from both PEM:: and PEM2K::.
Results of the test of NRT data transfers from the rehost computer were
successful for individual files which exist on CDHFB::.
PEM has been working with GSFC on allowing commanding of the PEM
instrument by creating DAPs on the PEM2K machine and submitting them to the
CMS. PEM2K:: now is able to communicate with the CMS, but the CMS does
not recognize PEM2K:: as an authorized site for instrument daily activity
plan submission. PEM is waiting for a solution from GSFC.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in a UARS Battery Status Review
Meeting on 10 May 2001. Daily processing of instrument
Daily Activity Plans continues with no problems. The
instrument operations schedule is now complete
to 28 June 2001.
***********
CDHF
***********
STATUS
- The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
***LEVEL0 for UARS DAYS 3514 and 3515 re-ingested with
additional data on current system.
2 days ATT V7.3 (PROD) (3514, 3515) (2R-2C=0L)
UARS CDHF REHOST:
- NRT user access testing occurring this week.
- LEVEL0 for UARS DAYS 3514 and 3515 was re-ingested with additional
data. Attitude and Observatory log were reprocessed for these days.
FDF indicated no reprocessing of Sensor software was necessary.
DAAC DATA:
- CLAES (OUTPUT 9) data created in May, 1999 has been loaded into the
DAAC Transfer Tables. DAAC has been notified.
- CLAES (OUTPUT 9) data created in June, 1999 has been previewed in
preparation for sending to the DAAC. The list of June files has been
forwarded to CLAES for review and approval.
***********
FOT
***********
REPORTING PERIOD
This report covers 28 April 2001 (Orbit 52635, GMT Day 118) through 04
May 2001 (Orbit 52740, GMT Day 124).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
AXIS 1&2 120/0102 52666 Power ON
SOLSTICE 122/0142 52696 SOLSTICE REBOOT
HALOE 123/1310 52719 Power OFF (TBAT TEMP)
The beta angle ranged between the angles of 36.1 degrees to 60.9 degrees
for this report period. The beta angle is now increasing towards a
maximum angle of 74 degrees on 09 May 2001 (DOY 129).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.8 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.2 and 1.3 degrees C. Additional
battery performance data for each battery on 27 April 2001 and 4 May
2001 are provided in Appendix D.
The clock error ranged from Ð10.8 to 10.5 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and
Out-Of-Limits occurrences are listed in Appendix B.
¥ OUT-OF-LIMITS = 04
¥ ALERTS = 02
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
¥ FOT = 25
¥ HALOE = 01
¥ AXIS = 01
¥ SOLSTICE = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
¥ FOT = 03
Three (3) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/119 - 2287.499116
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 01/067 show the observatory data loss to
be 300 hours, 18 minutes, 29.768 seconds. This shows no change since the
last report period. This is a 0.39 percent data loss, which equals a
99.61 percent data capture for the mission.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The new flight software patch uplinked on 25 April 2001 continues to be
monitored this week. The FOT is pleased with its overall performance
and is fine-tuning the TAM alignment by loading new alignment matrices.
An accurate alignment is needed to get ready for the loss of sun in the
field of view of the FSS on 5 May 2001.
MEETINGS
The FOT attended a UARS Monthly Status Meeting on Wednesday, May 02,
2001.The FOT attended a UARS FSW CCB Meeting on Thursday, May 03, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 30 May 2001 (DOY 150). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 04 May 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
52639 118/0434 01-270 NBTR NBTR PBK / RTS-58
52640 118/0653 01-270 NBTR NOOPS D-118
52647 118/1958 01-271 NBTR NOOPS D-119
52647 118/1958 01-272 OBC/CLK Clk Adj 20.6 to 0.00 msec/day
52653 119/0436 01-271 NBTR NBTR PBK / RTS-58
52655 119/0655 01-271 NBTR NOOPS D-119
52662 119/1818 01-273 NBTR NOOPS D-120
52666 120/0102 01-274 AXIS 1&2 Power ON
52668 120/0437 01-273 NBTR NBTR PBK / RTS-58
52670 120/0657 01-273 NBTR NOOPS D-120
52675 120/1452 01-275 MACS TAM ALIGNMENT
52677 120/1819 01-276 NBTR NOOPS D-121
52684 121/0620 01-276 NBTR NBTR PBK / RTS-58
52685 121/0659 01-276 NBTR NOOPS D-121
52691 121/1639 01-277 OBC Clk Adj 0.00 to 20.6 msec/day
52693 121/2042 01-279 NBTR NOOPS D-122
52696 122/0142 01-280 SOLSTICE SOLSTICE REBOOT
52696 122/0255 01-281 SOLSTICE SOLSTICE IGNORE CMD
52698 122/0439 01-279 NBTR NBTR PBK / NBTR REC CMDS
52702 122/1031 01-279 NBTR NBTR PBK / RTS-58
52703 122/1307 01-279 NBTR NOOPS D-122
52709 122/2140 01-283 NBTR NOOPS D-123
52713 123/0440 01-283 NBTR NBTR PBK / RTS-58
52718 123/1203 01-283 NBTR NOOPS D123
52719 123/1310 01-284 HALOE Power OFF (TBAT TEMP)
52721 123/1721 01-285 OBC Clk Adj fr 20.6 to 0.00 msec/day
52729 124/0519 01-287 SSPP enable SSPP FDC
52729 124/0621 01-288 SSPP SSPP FDC recovery & noops
52738 124/2001 01-289 NBTR PLAYBACK
________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 04 May 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
SOSNTEVLTG R-LO SOLSTICE N TUBE VOLTAGE 52696,7
SOSFTEVLTG R-LO SOLSTICE F TUBE VOLTAGE 52696,7
SOSGTEVLTG R-LO SOLSTICE G TUBE VOLTAGE 52696,7
HATBAT Y-HI HALOE BOLOMETER AREA TEMP 52715,6
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
ACHSKPG3REDL ACRIM RED LOW VOLTAGE 52709
________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 04 May 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
52647 118/2008 Clk Adj from 20.6 to 0.00 msec/day
52691 121/1650 Clk Adj from 0.00 to 20.6 msec/day
52721 123/1721 Clk Adj from 20.6 to 0.00 msec/day
________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 04 May 2001
27 Apr 2001 - GMT Day 117 Beta = 31.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 12.0 0.000 0.0
2 +11.2/-89.6 +6.11/+4.51 +28.8/-15.2 24.2 1.030 21.6
3 +22.4/-50.4 +5.15/+3.55 +28.8/-16.0 24.2 1.020 21.6
4 May 2001 - GMT Day 124 Beta = 60.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/+0.0 11.8 0.000 0.0
2 +00.0/-33.6 +3.87/+2.60 +22.0/-15.2 26.4 1.060 16.8
3 +22.4/-22.4 +2.92/+1.66 +21.2/-15.2 26.4 1.050 16.4
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
---- ---- ---------
-no changes V/T 5
________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 04 May 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-021 SOLSTICE 52696 122/0142 N/A Det. OFF and Tubes volt. low
01-022 ACRIM 52709 122/2140 N/A Red Low voltage alert
01-023 FOT 52728 124/0441 N/A SSPP FDC tripped
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-023 FOT 52728 124/0441 N/A SSPP FDC tripped.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-019 FOT see list Apr 01 14170 Generic Late Acquisition
01-021 SOLSTICE 52696 122/0142 N/A Det. OFF and Tubes volt. low
01-022 ACRIM 52709 122/2140 N/A Red Low voltage alert
SEVNTFW01/16